OpenStudy (anonymous):

medal included!! Which ratio is not equivalent to 45/120 ? A. 3/8 B. 5/12 C. 9/24 D. 90/240

OpenStudy (anonymous):

Are you in 7th grade? Maybe I can help u with this!

OpenStudy (anonymous):

yea

OpenStudy (skullpatrol):

What does "equivalent ratio" mean to you?

OpenStudy (anonymous):

good

OpenStudy (anonymous):

like somthing that equals the same

OpenStudy (skullpatrol):

What does the question ask for?

OpenStudy (anonymous):

oh is it d

OpenStudy (skullpatrol):

yes you are right, equivalent ratios are ratios that represent the SAME number.

OpenStudy (skullpatrol):

D But 90/240 = 45/120

OpenStudy (anonymous):

Sorry to was doing !y homework

OpenStudy (skullpatrol):

Which ratio is NOT equivalent to 45/120 ?

OpenStudy (anonymous):

A @skullpatrol

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Catswillonedayrule

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

can someone just help me

OpenStudy (michele_laino):

@coolkilla03 please note that the greatest common divisor between 45 and 120, is 15, so yuo have to divide both numerator and denominator by 15, try please

OpenStudy (michele_laino):

for example 45/15 = 3, noe do the same with the denominator, please, write your answer

OpenStudy (anonymous):

oh its 3/8

OpenStudy (michele_laino):

@coolkilla03 that's right!

OpenStudy (anonymous):

so it A

OpenStudy (michele_laino):

@coolkilla03 that's right!

OpenStudy (michele_laino):

sorry, please I have made an error, you answer is B. @coolkilla03

OpenStudy (michele_laino):

because if you multiply bot numerator and denominator of your ratio, by 2you will get: 90/120, namely D. whereas if you divide both numerator and denominator of your ratio by 5, you will get 9/24, namely C.

OpenStudy (michele_laino):

finally if you divide or multiply both numerator and denominator of your ratio by the same number, you will not never get 5/12, namely B.